Please help me solve this problem.
Given: f(x)=x^2+3x+1 g(x)= x^3+1 h(x)=x^2-2 d(x)= x+1

find (f o g)(1/3)

Answer by stanbon(48535)

f(x)=x^2+3x+1 g(x)= x^3+1 h(x)=x^2-2 d(x)= x+1
find (f o g)(1/3)
---------------------------
f[g(1/3)] = f[(1/3)^3 + 1]
= f[28/27]
= (28/27)^2 + 3(28/27) + 1
= 3781/729 or 5.186565
=====================
Cheers,
Stan H.